# Hybrid Latent Reasoning via Reinforcement Learning

### v1

@@ 5 Conclusion
In this work, we propose hybrid reasoning policy optimization (HRPO), a novel reinforcement learning (RL) framework that unifies discrete token sampling with continuous latent representations through a learnable gating mechanism. By gradually incorporating hidden features into sampled token embeddings, HRPO incentivizes LLMs to refine their reasoning strategies hybridly. Extensive evaluations on knowledge and STEM benchmarks demonstrate that HRPO outperforms both SFT and RL baselines, achieving consistent gains across diverse scenarios. Moreover, our analysis reveals that HRPO not only ensures stable hybrid latent reasoning but also triggers intriguing reasoning patterns, showing the HRPO’s potential and providing insights for further work in latent reasoning.
@@ Appendix A Implementation
For hybrid latent reasoning, our plug-and-play component is by design compatible with any LLM architecture. We initialize its linear layers with a uniform distribution from $[-1/\sqrt{|H|},1/\sqrt{|H|}]$, where $|H|$ denotes the hidden state dimension. The gating parameter $\Lambda$ is selected such that the quantity $a^{c}=\texttt{exp}(-c\cdot\texttt{softplus}(\Lambda))$ is drawn uniformly from $[r_{\mathrm{min}},0.999]$, with the scalar constant fixed at $c=8$ [5]. Tuning $r_{\mathrm{min}}$ adjusts the initial fraction of hidden states involved in hybrid reasoning; a larger value increases the proportion of sampled token embeddings and can be helpful for enhancing generation quality during the initial training phase. Similarly, the temperature hyperparameter $\tau$ in Equation 3 can be tuned for optimal task performance, although HRPO remains robust across a wide range of $\tau$ values. To efficiently train the LLMs with HRPO, we patch the models with optimized kernel implementations [^2] https://github.com/unslothai/unsloth and employ low-rank adaptation (LoRA) [15]. The default choice of hyperparameters are reported in Table 4 for HRPO experiments.
@@
Learning Rate (Linear in Equation 4) 1e-4
Learning Rate ($\Lambda$ in Equation 4) 1e-3
@@
The hyperparameters are selected empirically to balance efficiency and performance, and thanks to HRPO’s lightweight design and additional optimizations, our framework can run on a single GPU across diverse tasks. Additionally, we apply a larger weight-decay coefficient to (1) enhance HRPO training stability and (2) encourage the gating towards incorporating more latent representations (since smaller positive $\Lambda$ values increase the hidden ratio $\sqrt{1-a_{t}^{2}}$). For simpler knowledge tasks and GSM8k, we fix the HRPO group size at 4, which already delivers strong performance. For more challenging benchmarks, namely MATH, MATH500, MMLU-ST and ARC-C, we instead generate 8 hybrid completions for each query. As for prompt and completion lengths, we select them empirically based on our observations, and the selected values are summarized in Table 5.
@@
For both training and evaluation, we build each prompt by prepending a system message that directs the LLM to perform step-by-step internal reasoning before generating its final answer. The user query is then appended, and the entire input is formatted with the model chat template. Different from prior work [10, 18], we adopt the minimalist delimiter #### to separate the model’s hybrid reasoning span from its final answer. This is because the delimiter tokenizes as a single unit, adding no length overhead while providing a clear signal to switch from hybrid latent reasoning to autoregressive answer generation. We also penalize repeated occurrences of the delimiter within the completion (by assigning 0 reward regardless answer correctness) to prevent the model from early termination of hybrid reasoning. We illustrate full prompts for different type of tasks, showing the system message and example queries in Figure 7, Figure 8 and Figure 9, respectively.
@@
For each question in our knowledge-intensive QA setup, we embed the query with E5 embedding model [42]. The entire English Wikipedia 2020 dump is pre-encoded with the same model, after which we perform approximate nearest neighbor (ANN) search and select the three highest-scoring documents. These top-3 passages are concatenated to form the external context fed to the LLM, as illustrated in Figure 7. In our evaluation, we generate tokens using greedy decoding and compute latent representations according to Equation 3, thereby ensuring the reproducibility of our results. For outcome-based reward and evaluation settings on knowledge tasks, we report exact match scores on val / test splits following [50, 51, 18]. For mathematical (GSM8k, MATH and MATH500) and multiple-choice datasets (MMLU-ST and ARC-C), we follow [49] for post-processing and scoring.
@@ Appendix B Additional Results
Comparison to Latent Reasoning Methods. In addition to strong RL methods such as PPO and GRPO in our main experiments, we also benchmark the proposed HRPO against additional latent reasoning baselines. Specifically, we evaluate HRPO, Coconut and CODI on the GSM8K and MATH reasoning datasets, all using the 1.5B Qwen backbone. For Coconut, we train with its augmented CoT data (no MATH split is available), whereas for CODI we adopt the original datasets’ CoT trajectories. The results are reported in Table 6. We observe: (1) HRPO achieves the best accuracy on both datasets, with 9.42% and 23.63% respective gains over the best performing latent reasoning baseline CODI. (2) Even compared to distilled CoT from a significantly larger model QwQ, HRPO still scores consistent improvements on both datasets, showing the effectiveness of our hybrid latent reasoning. (3) Coconut lags behind on GSM8k, indicating limitations of latent reasoning by compressing CoT tokens, whereas CODI improves substantially with CoT SFT but still trails Distilled CoT and HRPO. Overall, HRPO achieves the best performance against all baselines, demonstrating its consistent advantages over CoT distillation and prior latent reasoning methods.
@@
Sensitivity Analysis for $\Lambda$ and $\tau$. In addition to the results reported in Table 3, we further present the performance of various $\Lambda$ initializations on the Qwen 3B model, as shown in Table 7. Our observations echo the same trends on the 1.5B backbone: a smaller initial $r_{\mathrm{min}}$ consistently benefits both knowledge and STEM tasks. Moreover, performance peaks when $r_{\mathrm{min}}$ is selected either lower or higher, and drops slightly within the intermediate range of $[0.98,0.999]$. We also examine the sensitivity of the $\tau$ hyperparameter used to construct the interpolated embeddings and present the corresponding results for both backbone models in Table 8. The training rewards and completion lengths for GSM8k, MATH and the knowledge tasks are shown in Figure 10, Figure 11 and Figure 12. We note that choosing $\tau$ in the range of 0.5 – 0.7 offers a reliable balance of efficiency and accuracy, as the performance often peaks around this interval for both backbone models. Overall, we find that HRPO benefits from a smaller initial $r_{\mathrm{min}}$, which outperforms larger $r_{\mathrm{min}}$ settings and highlights the value of latent representations for complex reasoning. Moreover, HRPO is robust to the choice of $\tau$, where the performance scores remain stable with only minor fluctuations at the extremes.
@@
Model Qwen 1.5B Qwen 3B
@@
Additional Analysis for $\Lambda$ Initialization. We further provide an expanded analysis of how varying $r_{\mathrm{min}}$ in the initialization of $\Lambda$ affects training dynamics with the larger Qwen 3B backbone. Figures Figure 13, Figure 14, Figure 15 and Figure 16 plot the reward and completion length curves for the knowledge tasks, GSM8k, MATH and MMLU-ST / ARC-C respectively. Overall, our findings here echo the observations in Section 4.3: different $r_{\mathrm{min}}$ values exhibit similarly high training stability and preserve the LLM’s generative capabilities, but selecting a smaller $r_{\mathrm{min}}$ (i.e., a larger initial hidden ratio) generally accelerates convergence and shortens generated completions. Nevertheless, these benefits are less pronounced for the 3B backbone than for the 1.5B counterpart, which we attribute to the fewer update steps and trainable parameters in HRPO. In summary, our analysis shows that HRPO preserves stable training dynamics and effectively leverages LLMs’ intrinsic reasoning patterns across $r_{\mathrm{min}}$ values; moreover, choosing a smaller $r_{\mathrm{min}}$ further enhances convergence and yields shorter generated sequences, which can be especially beneficial for smaller-scale LLMs.
@@ Appendix C Qualitative Analysis
To further highlight HRPO’s reasoning patterns, we present additional qualitative examples. Each example provides the reasoning trace by decoding the sampled tokens from the hybrid reasoning process, and we include both successful and erroneous cases across different tasks in the following. The correct examples are provided in Figure 17, Figure 18, Figure 19, Figure 20, Figure 21, where as the mistakes are provided in Figure 22, Figure 23, Figure 24, Figure 25, Figure 26, we show the raw strings and omit the options / contexts in the examples due to space constraints.
@@
As for reasoning errors, we also identify several common patterns: (1) Cross-lingual mistakes arising from limited parametric or contextual knowledge, as in Figure 22 and Figure 23. (2) Correct answers that violate the predefined format and thus receive a zero score (Figure 24). (3) Repetitive loops that continue until the response hits the maximum completion length (Figure 25). (4) Cross-lingual reasoning that is nonetheless truncated by the length limit (Figure 26). Overall, these patterns indicate that, while HRPO effectively integrates discrete and latent representations in its internal reasoning process, it may be further enhanced through refined output formatting (e.g., potentially with a format reward), extended optimization schedules with conservative learning, increased model parameters, and longer context / generation capabilities, pointing to promising directions for future research.

### v2

@@ 5 Conclusion
In this work, we propose hybrid reasoning policy optimization (HRPO), a novel reinforcement learning (RL) framework that unifies discrete token sampling with continuous latent representations through a learnable gating mechanism. By gradually incorporating hidden features into sampled token embeddings, HRPO incentivizes LLMs to refine their reasoning strategies hybridly. Extensive evaluations on knowledge and STEM benchmarks demonstrate that HRPO outperforms both SFT and RL baselines, achieving consistent gains across diverse scenarios. Moreover, our analysis reveals that HRPO not only ensures stable hybrid latent reasoning but also triggers intriguing reasoning patterns, showing its potential in reasoning-intensive settings and providing insights for RL-based continuous space learning. While promising, we recognize that HRPO introduces additional computation overhead, the on-policy design may reduce large-scale training efficiency, and its continuous representations can be less transparent. Therefore, future work will aim to address these limitations by exploring simpler designs, off-policy extensions and advanced latent reasoning techniques to improve both the interpretability and efficiency of HRPO.
@@ Appendix A Implementation
For hybrid latent reasoning, our plug-and-play component is by design compatible with any LLM architecture. We initialize its linear layers with a uniform distribution from $1/\sqrt{|H|},1/\sqrt{|H|}]$, where $|H|$ denotes the hidden state dimension. The gating parameter $\Lambda$ is selected such that the quantity $a^{c}=\texttt{exp}(-c\cdot\texttt{softplus}(\Lambda))$ is drawn uniformly from $r_{\mathrm{min}},0.999]$, with the scalar constant fixed at $c=8$ [5]. Tuning $r_{\mathrm{min}}$ adjusts the initial fraction of hidden states involved in hybrid reasoning; a larger value increases the proportion of sampled token embeddings and can be helpful for enhancing generation quality during the initial training phase. Similarly, the temperature hyperparameter $\tau$ in Equation˜3 can be tuned for optimal task performance, although HRPO remains robust across a wide range of $\tau$ values. To efficiently train the LLMs with HRPO, we patch the models with optimized kernel implementations [^3] https://github.com/unslothai/unsloth and employ low-rank adaptation (LoRA) [15]. The default choice of hyperparameters are reported in Table˜4 for HRPO experiments.
@@
Learning Rate (Linear in Equation˜4) 1e-4
Learning Rate ($\Lambda$ in Equation˜4) 1e-3
@@
The hyperparameters are selected empirically to balance efficiency and performance, and thanks to HRPO’s lightweight design and additional optimizations, our framework can run on a single GPU across diverse tasks. Additionally, we apply a larger weight-decay coefficient to (1) enhance HRPO training stability and (2) encourage the gating towards incorporating more latent representations (since smaller positive $\Lambda$ values increase the hidden ratio $\sqrt{1-a_{t}^{2}}$). For simpler knowledge tasks and GSM8k, we fix the HRPO group size at 4, which already delivers strong performance. For more challenging benchmarks, namely MATH, MATH500, MMLU-ST and ARC-C, we instead generate 8 hybrid completions for each query. As for prompt and completion lengths, we select them empirically based on our observations, and the selected values are summarized in Table˜5.
@@
For both training and evaluation, we build each prompt by prepending a system message that directs the LLM to perform step-by-step internal reasoning before generating its final answer. The user query is then appended, and the entire input is formatted with the model chat template. Different from prior work [10, 18], we adopt the minimalist delimiter #### to separate the model’s hybrid reasoning span from its final answer. This is because the delimiter tokenizes as a single unit, adding no length overhead while providing a clear signal to switch from hybrid latent reasoning to autoregressive answer generation. We also penalize repeated occurrences of the delimiter within the completion (by assigning 0 reward regardless answer correctness) to prevent the model from early termination of hybrid reasoning. We illustrate full prompts for different type of tasks, showing the system message and example queries in Figure˜7, Figure˜8 and Figure˜9, respectively.
@@
For each question in our knowledge-intensive QA setup, we embed the query with E5 embedding model [42]. The entire English Wikipedia 2020 dump is pre-encoded with the same model, after which we perform approximate nearest neighbor (ANN) search and select the three highest-scoring documents. These top-3 passages are concatenated to form the external context fed to the LLM, as illustrated in Figure˜7. In our evaluation, we generate tokens using greedy decoding and compute latent representations according to Equation˜3, thereby ensuring the reproducibility of our results. For outcome-based reward and evaluation settings on knowledge tasks, we report exact match scores on val / test splits following [50, 51, 18]. For mathematical (GSM8k, MATH and MATH500) and multiple-choice datasets (MMLU-ST and ARC-C), we follow [49] for post-processing and scoring.
@@ Appendix B Additional Results
Comparison to Latent Reasoning Methods. In addition to strong RL methods such as PPO and GRPO in our main experiments, we also benchmark the proposed HRPO against additional latent reasoning baselines. Specifically, we evaluate HRPO, Coconut and CODI on the GSM8K and MATH reasoning datasets, all using the 1.5B Qwen backbone. For Coconut, we train with its augmented CoT data (no MATH split is available), whereas for CODI we adopt the original datasets’ CoT trajectories. The results are reported in Table˜6. We observe: (1) HRPO achieves the best accuracy on both datasets, with 9.42% and 23.63% respective gains over the best performing latent reasoning baseline CODI. (2) Even compared to distilled CoT from a significantly larger model QwQ, HRPO still scores consistent improvements on both datasets, showing the effectiveness of our hybrid latent reasoning. (3) Coconut lags behind on GSM8k, indicating limitations of latent reasoning by compressing CoT tokens, whereas CODI improves substantially with CoT SFT but still trails Distilled CoT and HRPO. Overall, HRPO achieves the best performance against all baselines, demonstrating its consistent advantages over CoT distillation and prior latent reasoning methods.
@@
Sensitivity Analysis for $\Lambda$ and $\tau$. In addition to the results reported in Table˜3, we further present the performance of various $\Lambda$ initializations on the Qwen 3B model, as shown in Table˜7. Our observations echo the same trends on the 1.5B backbone: a smaller initial $r_{\mathrm{min}}$ consistently benefits both knowledge and STEM tasks. Moreover, performance peaks when $r_{\mathrm{min}}$ is selected either lower or higher, and drops slightly within the intermediate range of $[0.98,0.999]$. We also examine the sensitivity of the $\tau$ hyperparameter used to construct the interpolated embeddings and present the corresponding results for both backbone models in Table˜8. The training rewards and completion lengths for GSM8k, MATH and the knowledge tasks are shown in Figure˜10, Figure˜11 and Figure˜12. We note that choosing $\tau$ in the range of 0.5 – 0.7 offers a reliable balance of efficiency and accuracy, as the performance often peaks around this interval for both backbone models. Overall, we find that HRPO benefits from a smaller initial $r_{\mathrm{min}}$, which outperforms larger $r_{\mathrm{min}}$ settings and highlights the value of latent representations for complex reasoning. Moreover, HRPO is robust to the choice of $\tau$, where the performance scores remain stable with only minor fluctuations at the extremes.
@@
Model Qwen2.5-1.5B Qwen2.5-3B
@@
Additional Analysis for $\Lambda$ Initialization. We further provide an expanded analysis of how varying $r_{\mathrm{min}}$ in the initialization of $\Lambda$ affects training dynamics with the larger Qwen 3B backbone. Figures Figure˜13, Figure˜14, Figure˜15 and Figure˜16 plot the reward and completion length curves for the knowledge tasks, GSM8k, MATH and MMLU-ST / ARC-C respectively. Overall, our findings here echo the observations in Section˜4.3: different $r_{\mathrm{min}}$ values exhibit similarly high training stability and preserve the LLM’s generative capabilities, but selecting a smaller $r_{\mathrm{min}}$ (i.e., a larger initial hidden ratio) generally accelerates convergence and shortens generated completions. Nevertheless, these benefits are less pronounced for the 3B backbone than for the 1.5B counterpart, which we attribute to the fewer update steps and trainable parameters in HRPO. In summary, our analysis shows that HRPO preserves stable training dynamics and effectively leverages LLMs’ intrinsic reasoning patterns across $r_{\mathrm{min}}$ values; moreover, choosing a smaller $r_{\mathrm{min}}$ further enhances convergence and yields shorter generated sequences, which can be especially beneficial for smaller-scale LLMs.
Statistical Significance Analysis on the Improvements of HRPO. In our main experiments, we follow the standard practice of using greedy decoding for pass@1 evaluation, ensuring our results are easy to evaluate and reproducible. To evaluate the significance of the performance gains of HRPO, we conduct additional sampling-based evaluations on the STEM tasks, which exhibit greater variance compared to greedy decoding. Averaged results are presented in Table˜9, with statistically significant outcomes (paired t-test, $p<0.05$) highlighted in bold. These results show that HRPO consistently outperforms PPO and GRPO across both backbones on all benchmark datasets. For the 1.5B backbone, t-tests confirm these gains are statistically significant in three out of five tasks. The improvements are even more pronounced with the 3B model, which achieves an average gain of +1.4% and shows statistical significance in four out of five comparisons. These findings demonstrate that our hybrid-RL framework, HRPO, not only delivers reliable performance gains over established baselines but also does so with high statistical confidence across the majority of STEM tasks.
Table 9: Significance test on HRPO’s performance improvements.
Qwen2.5-1.5B
GSM8k MATH MATH500 MMLU-ST ARC-C
PPO 0.701 0.505 0.511 0.551 0.716
GRPO 0.710 0.510 0.512 0.554 0.722
HRPO 0.712 0.515 0.517 0.565 0.731
Qwen2.5-3B
GSM8k MATH MATH500 MMLU-ST ARC-C
PPO 0.825 0.597 0.600 0.574 0.802
GRPO 0.827 0.595 0.599 0.577 0.808
HRPO 0.838 0.606 0.609 0.585 0.815
@@ Appendix C Qualitative Analysis
To further highlight HRPO’s reasoning patterns, we present additional qualitative examples. Each example provides the reasoning trace by decoding the sampled tokens from the hybrid reasoning process, and we include both successful and erroneous cases across different tasks in the following. The correct examples are provided in Figure˜17, Figure˜18, Figure˜19, Figure˜20, Figure˜21, where as the mistakes are provided in Figure˜22, Figure˜23, Figure˜24, Figure˜25, Figure˜26, we show the raw strings and omit the options / contexts in the examples due to space constraints.
@@
As for reasoning errors, we also identify several common patterns: (1) Cross-lingual mistakes arising from limited parametric or contextual knowledge, as in Figure˜22 and Figure˜23. (2) Correct answers that violate the predefined format and thus receive a zero score (Figure˜24). (3) Repetitive loops that continue until the response hits the maximum completion length (Figure˜25). (4) Cross-lingual reasoning that is nonetheless truncated by the length limit (Figure˜26). Overall, these patterns indicate that, while HRPO effectively integrates discrete and latent representations in its internal reasoning process, it may be further enhanced through refined output formatting (e.g., potentially with a format reward), extended optimization schedules with conservative learning, increased model parameters, and longer context / generation capabilities, pointing to promising directions for future research.